sue is now 7 times as old as louis. in 6 yearss, sue will be 5 times as old a louis. can you help me find the present age for both sue and louis
Answer by Earlsdon(6294)

Try this:
Let S = Sue's present age and L = Louis's present age.
1) S = 7L "Sue is 7 times as old as Louis."
2) S+6 = 5(L+6) "In 6 years, Sue will be 5 times as old as Louis (will be in 6 years time)" Substitute equation 1) into equation 2) and solve for L.
2a) (7L)+6 = 5(L+6) Simplify.
7L+6 = 5L+30 Subtract 5L from both sides.
2L+6 = 30 Subtract 6 from both sides.
2L = 24 Divide both sides by 2.
L = 12 and...
S = 7L
S = 7(12)
S = 84